3. AN OBJECT STARTS FROM REST AND UNIFORMLY ACCELERATION AT A RATE OF 2m/sec2 FOR 5 SECOND.

A. WHAT IS THE DISPLACEMENT DURING THIS 5 SECOND TIME PERIOD?

B. WHAT IS THE OBJECT FINAL VELOSITY?


1
Expert's answer
2021-06-01T10:50:05-0400

a)


"s=0.5at^2\\s=0.5(2)(5)^2=25\ m"

b)


"v=at\\v=2(5)=10\frac{m}{s}"